A Glass beaker is filled to the top with 800 cm3 of mercury. if the temperature of the mercury is raised from 0C to 368 K, how much mercury will spill out?
1
Expert's answer
2011-02-23T07:07:28-0500
The thermal expansion of mercury is given by V(t) = V(1 + 1.82 x 10-4t + 7.8 x 10-9t2), where t is in °C, and V is the volume at 0°C.
Thus V = 800 (1 + 1.82 x 10-4*368 + 7.8 x 10-93682) = 854.42 cm3.
854.42 - 800 = 54.42 cm3 will spill out of the glass.